# Transaction Profiling and Address Role Inference in Tokenized U.S. Treasuries

### v1

@@ 5 Address Role Predictive Modeling
In this section, we focus on address role predictive modeling within the RWA transactions, with the goal of inferring the financial roles of addresses, i.e., not their identity per se, but the economic or operational behaviour they instantiate on-chain, such as traders, execution bots, treasury managers protocol contracts. Even at this level of approximation, functional role inference enables us to potentially contextualize address behaviour, differentiate structural actors in token flows, and support future analysis of token circulation, institutional activity.
Such financial roles can, in part, be derived from existing address labels where available from online resources. However, Labeling on-chain addresses remains a manual and epistemically incomplete task, currently reliant on community-curated, e.g., Dune community-based crypto analytics  [^4] https://dune.com/, or industry-maintained ontologies, e.g., Arkham  [^5] https://intel.arkm.com/, Ethplorer  [^6] https://ethplorer.io/zh/ that emerge retroactively through crowd-sourced investigation, custodial disclosures, or industrial exogenous tagging. These labeling processes are temporally cumulative, i.e., they evolve incrementally over time, contingent on the visibility and social salience of individual addresses. As a result, there exists a persistent gap in functional address understanding at the time of analysis, which necessitates predictive modeling frameworks capable of approximating behavioural roles with minimal supervision. Such automation and approximation provide a practical reference for analysts, partially reducing the cost for manual investigation by offering behaviourally grounded priors that can guide interpretation for manual labeling.
@@
Concretely, we advocate for the inference of financial behaviourally grounded roles: financial controllers treasuries, execution (arbitrage) bots, and retail traders, as proxy labels. In this schema, treasuries refer to multisignature-controlled addresses or custodial vaults (e.g., Gnosis Safe contracts, DAO-managed wallets) that act as long-term capital reserves, execute protocol expenditures, or manage liquidity across chains. The category of execution bots encompasses addresses engaged in high-frequency, low-latency, and often adversarial strategies such as sandwich attacks, MEV extraction, flashloan arbitrage, and Flashbots relaying, whose behavioural signatures deviate from standard market participation and suggest automation and profit maximization. In contrast, retail traders include externally owned accounts (EOAs) that interact primarily with DEXs, NFT marketplaces, or aggregator routers. Such modeling serves both to reduce annotation overhead and potentially enable functional interpretability in transaction graph-level analysis.
5.1 Samples and labels
@@
We aggregate transactions from three representative tokenized real-world assets (RWAs): USDY (Ethereum), BENJI (excluding stellar), and BUIDL. These RWS U.S treasury tokens were selected due to their active cross-chain circulation, institutional provenance, and higher visibility and density of human-readable annotations within community-maintained labeling available sources, making them suitable for semi-supervised role inference.
@@
We extract the naming tags for each address using Dune SQL from the metadata of Dune labels.addresses table  [^7] https://dune.com/data/labels.addresses: an open, curated repository of community-submitted and platform-extracted multi-chain address labels. Each entry includes fields such as name, address, blockchian, source, contributor, etc. We use the name field in the table, which is the naming tags, then apply a set of regular-expression-based rules over the name field to assign coarse-grained functional roles. These rules approximate the aforementioned financial behavioural classes. Table 5 summarizes the pattern-matching logic used to infer each role:
@@
Table 5: Regular expression rules used for coarse address role labeling.
@@
In total, the sampled dataset contains 10,055 transactions between 2023-09-18 and 2025-04-23, covering 815 unique addresses observed as senders or recipients. Using this labeling scheme, we label the address set into four classes: 520 Trader, 33 Bot, 44 Treasury, and 218 Other. These labels serve as training targets for our predictive modeling pipeline, allowing us to benchmark role inference performance under weak supervision.
5.2 Models
RWA transaction graphs exhibit latent financial hierarchy: central actor nodes (e.g., treasuries, issuers, custodians) initiate and coordinate flows, while peripheral nodes (e.g., traders, bots) interact sparsely, locally, or opportunistically with the core. The negative curvature of hyperbolic space provides a proper inductive bias for embedding such structure, enabling compact representations that naturally separate high-degree core nodes near the origin from low-degree boundary nodes along exponentially expanding geodesics. Hence, nodes residing at small hyperbolic radius embody high-hierarchy, systemically central nodes (close to coordinate origin) possibly multisig treasuries, protocol routers, and liquidity hubs whereas larger radius nodes reflect lower-hierarchy participants across the RWA transaction graph. A formal derivation showing that hyperbolic radius encodes hierarchical depth in tree-like graphs where the higher-hierarchy is closer to the originate is provided in Appendix 7, and an illustrative Figure 3 is presented.
Therefore, we propose a hyperbolic node-level representation learning method combined with a feedforward neural networks that integrates transactional features, metadata-driven features including Liquidity-to-Average Ratio (LAR), and hyperbolic (Poincaré) geometry embeddings. The pipeline first encodes latent hierarchy via Poincaré distance-based optimization, then augments each node with hierarchical depth statistics and optional topological features, which are processed by a neural network classifier to infer address roles.
Figure 3: Illustrative example of a rendering of a blockchain transaction graph and its corresponding Poincaré disk embedding. Nodes emulate a three-tier addresses: the red address anchors global liquidity, the green addresses relay funds and manage liquidity pools, and a ring of peripheral trader blue addresses engages sporadically with the core. Embedding the same adjacency structure in the negatively curved Poincaré model separates tiers by geodesic radius.
5.2.1 Poincaré Node Rrepresentation Leanring
Each data point in our framework corresponds to a node $v\in V$ within a token transaction graph $G=(V,E)$, where nodes represent blockchain addresses and directed edges $(u,v)\in E$ denote value transfers. The objective is to learn a high dimensional representation (embedding vector) $\mathbf{z}_{v}\in\mathbb{B}^{d}$ for each address $v$, where $\mathbb{B}^{d}$ is the $d$-dimensional Poincaré ball: a Riemannian manifold of constant negative curvature:
@@ 5.2.1 Poincaré Node Rrepresentation Leanring
[…]
@@
Each update step is followed by projection back onto the open ball to ensure validity:
@@
where $\gamma>0$ is a margin hyperparameter and $[\cdot]_{+}=\max(0,\cdot)$ denotes the hinge operator. To align learned radii with graph-theoretic centrality, we regularize embedding norms against normalized degree:
@@
where $\beta$ (default 0.1) controls the strength of radial regularization. This objective is minimized over certain epochs using intrinsic gradients computed in the Riemannian manifold, followed by Möbius updates and projection. After optimization, each node $v\in V$ is assigned a $\mathbf{z}_{v}\in\mathbb{B}^{64}$, a 64-dimensional hyperbolic embedding that encodes its topological position and connectivity within the latent hierarchy of the transaction graph.
Refinement via Liquidity-to-Average Ratio (LAR) While Poincaré embeddings encode topological position in the latent transactional hierarchy, other features related to token flow transferred within transactions, i.e., the temporal liquidity behaviour that distinguishes roles such as passive holders, programmatic bots, or asymmetric treasuries. To incorporate such information, we define the following over nodes computed on transaction flows.
Definition 5.1 (Liquidity-to-Average Ratio (LAR)).
Let $(u,v)\in E$ denote a directed transaction edge, and let $\mu_{uv}$ and $\sigma_{uv}$ denote the mean and standard deviation of transfer values from $u$ to $v$ within a time window $[t,t+\Delta]$. Let $\texttt{in}(v)$ and $\texttt{out}(v)$ denote the total inflow and outflow of $v$ over the same interval. Then the Liquidity-to-Average Ratio (LAR) for $(u,v)$ is defined as:
@@
{observation}
Nodes with low liquidity and irregular behaviour are downweighted during smoothing, preserving geometric sharpness for high-confidence addresses (e.g., well-funded treasuries), while dampening noise from sparsely observed actors.
5.2.2 Hierarchical Radius-based Features
$r_{v}=2\cdot\tanh^{-1}(\|\mathbf{z}_{v}\|),$
which reflects the geodesic distance from the origin and serves as a continuous proxy for depth in the hierarchy, i.e., nodes closer to the origin are higher in the hierarchy, while those near the boundary are deeper.
To capture the local structure around node $v$, we define its $k$-hop neighbourhood $\mathcal{N}_{k}(v)\subseteq V$ (excluding $v$) and collect neighbour radii $\{r_{u}:u\in\mathcal{N}_{k}(v)\}$. From this, we compute an 11-dimensional hierarchical feature vector $\mathbf{h}_{v}\in\mathbb{R}^{11}$ comprising the following components:
$r_{v}$ Absolute depth (node’s self radius)
$\mathbf{b}_{v}$ 4-bin histogram of relative radii $(r_{u}-r_{v})$ over range $[-1,1]$
Altogether, the hierarchical feature vector is given by:
5.2.3 Neural Network Inferencer
After obtaining the Poincaré position embedding, and hierarchical radius-based features, we train a supervised neural classifier to infer coarse address roles. Each node $v\in V$ is represented by a concatenated input vectors:
  • $\mathbf{z}_{v}\in\mathbb{R}^{d}$ (constrained to $\mathbb{B}^{d}$): Poincaré embedding (cf. Section 5.2.1);
  • $\mathbf{h}_{v}\in\mathbb{R}^{11}$: hierarchical features computed from local radius-based statistics (cf. Section 5.2.2);
@@ 5.2.3 Neural Network Inferencer
5.2.4 Experiment Result
We evaluate our role classifier using a stratified train/test split with an 80/20 ratio over the set of labeled addresses with the four coarse-grained role classes (cf. Section 5.1). The model is optimized using the AdamW [13] optimizer with a learning rate of $10^{-3}$ and a weight decay of $0.2$. Training is conducted for up to 2k epochs with early stopping based on macro-F1 score on the validation set, using a patience threshold of 10 epochs.
The best-performing model checkpoint (by macro-F1) is selected for final evaluation. We conduct ablation study using with the input features of Poincaré embeddings $\mathbf{z}_{v}\in\mathbb{R}^{64}$, hierarchical descriptors $\mathbf{h}_{v}\in\mathbb{R}^{11}$ ($k=1$), and topology-aware node embeddings $\mathbf{r}_{v}\in\mathbb{R}^{64}$, with feature subsets in combination under different ablation settings.
We named our model PoincaVec and compare the method against three established node representation baselines: Node2Vec [10], Role2Vec [1], and FeatherNode [19]. We select Role2Vec, and FeatherNode as baselines for comparison due to their established efficacy in modeling blockchain transaction graphs, as evidenced in prior evaluations [8, 6]. We adopt the same hyperparameter settings used in prior blockchain transaction address representation learning works [17, 6] , setting the context size $10$, embedding dimension $64$, walk length $5$, number of walks per node $10$ for both Node2Vec and Role2Vec, while FeatherNode is configured SVD iterations 20. The baseline Node2Vec [10] captures basic neighbour homophily-based proximity via biased random walks, while Role2Vec [1] models structural equivalence by clustering nodes based on topological statistics (e.g., degree, triangle count, k-core number) into roles (types) and learning role-level embeddings where the nodes share the same role obtain the same embedding vector. FeatherNode [19] incorporates spectral and node-level distributional features from attributed graphs by applying Fourier transforms of empirical distributions to neighbourhood feature aggregations. These aggregations are subsequently compressed using truncated Singular Value Decomposition (SVD), yielding node embeddings that are invariant to permutation and sensitive to local graph structure.
Table 6: Role classification performance on the RWA dataset. H: hierarchical radius-based features; T: topology-aware node features.
@@ 5.2.4 Experiment Result
Table 6 presents the classification performance of PoincaVec and several baseline embedding methods on the RWA dataset. Among the baselines, Node2Vec and Role2Vec provide modest F1 scores of 0.654 and 0.662, respectively, while FeatherNode underperforms 0.497, likely due to the high sparsity and low feature homogeneity in financial transaction graphs.
Without hierarchical or walk-based features, PoincaVec (w/o H, w/o T), which uses only Poincaré embeddings, outperforms all baselines with an F1 of 0.694. Concatenating hierarchical radius features (w/ H, w/o T) maintains competitive performance, while integrating topology-aware node vectors derived from truncated random walks (w/ H, w/ T) leads to a result F1 of 0.726. These results demonstrate that curvature-aware embeddings, when enriched with role-informed and structural signals, provide competitive representational capacity for role inference in blockchain transaction graphs.
Evaluation on External Blockchain Transaction Datasets To further evaluate the generalization capacity of our PoincaVec architecture beyond the RWA role classification task, we test it on multiple publicly available blockchain transaction graph datasets with labels. These datasets include addresses labeled as fraudulent or scam-related, whose topological irregularities render them particularly amenable to detection using curvature-aware embeddings and hierarchical structural descriptors. Given that fraud or anomalous addresses often lie structurally at the periphery or occupy non-homophilic positions in the graph, the curvature-aware inductive bias of PoincaVec may help reveal subtle hierarchical distinctions absent in flat Euclidean spaces.
Ethereum Transaction [31]
An Ethereum transaction transaction dataset constructed from historical blockchain records, containing 2,973,489 addresses and 13,551,303 edges. A total of 1,165 addresses are labeled as illicit based on curated forensic data, while the rest remain unlabeled. This dataset has an average node degree of 4.56.
AscendEXHacker [30]
A subgraph (network) extracted via Etherscan’s Heist label, focusing specifically on the 2021 AscendEX exchange exploit. It contains 6,642 addresses across 29,074 transactions. Ground-truth annotations label 84 addresses as direct heist participants, while 638 addresses are identified as DEXs or Uniswap-related service accounts.
PlusTokenPonzi [30]
A transaction dataset linked to the PlusToken Ponzi scheme, one of the largest crypto scams to date. It comprises 34,521 unique accounts and 58,049 transactions, of which 30,782 addresses are explicitly identified as scam participants.
Ethereum Classic Dataset [2]
A labeled transaction network from the Ethereum Classic (ETC) network, compiled using EtherscamDB scam reports. The dataset comprises 73,034 nodes and 71,250 edges, with 2,357 addresses labeled as scammers based on crowdsourced and externally verified annotations.
Table 7: Role classification performance across four blockchain transaction datasets.

### v2

@@ V Address Role Predictive Modeling
We focus on address role predictive modeling within the transactions, inferring (classifying) the financial roles of addresses, i.e., not the identity per se, but the financial or operational role. Financial role inference helps potentially contextualize address behaviour, differentiate structural actors in token flows, and support future analysis of token activity.
Such financial roles can, in part, be derived from existing address labels. However, on-chain addresses are overwhelmingly unlabeled. Inferring transaction intent, such as distinguishing flows from secondary-market trades, requires role inference from activity traces, which is nontrivial. Community-curated sources [15, 6, 18] and heuristics-based tagging in prior work [31, 49, 26] provide ex post labels. These labeling processes are evolving only as addresses gain visibility and social salience, and heuristic methods similarly rely on manually observed patterns, limiting their applicability to new transaction adoption. Graph-based machine learning models offer an improved approach by enabling automatic, ex ante role inference from activity traces, approximating financial roles directly from transactions that can be updated continuously [53, 22]. Such automation supplies behaviourally grounded priors that can guide interpretation, narrow the search space for manual investigation, to reduce the cost required for future comprehensive labeling.
@@
Concretely, we advocate for the inference of financial behaviour-grounded roles: financial controllers of treasury, execution (arbitrage) bots, and retail traders, as proxy labels. These classes map to distinct economic functions: treasury issuance or redemption liquidity, arbitrageurs, and retail participation, enabling cleaner flow separation and measurement of adoption. The capacity for such analysis has been widely needed across academic, industry, and policy communities (e.g., risk, compliance workflows, market monitoring) [30, 11, 30], motivating our focus on these three classes of roles. Treasuries refer to multisignature-controlled addresses or custodial vaults (e.g., Gnosis Safe contracts, DAO-managed wallets) that act as long-term capital reserves, execute protocol expenditures, or manage liquidity across chains. Execution bots engage in high-frequency, low-latency strategies (e.g., sandwiching, MEV extraction, flash-loan arbitrage, Flashbots relaying) whose signatures deviate from standard participation. Aggressive MEV bots can distort price discovery and congest networks, motivating explicit identification, as shown in prior studies [32, 25]. In contrast, retail traders include externally owned accounts (EOAs) that interact primarily with DEXs, NFT marketplaces, or aggregator routers.
V-A U.S. Treasury Tokens Dataset Annotation
@@
We aggregate transactions from three representative tokenized real-world assets (RWAs): USDY (Ethereum), BENJI (excluding Stellar), and BUIDL. These RWA U.S. Treasury tokens were chosen for their active cross-chain circulation, institutional provenance [29], and relatively more dense community-labeled annotations, making them well-suited for semi-supervised role inference.
@@
We extract the naming tags for each address using Dune SQL from the metadata of Dune labels.addresses [16]: a curated repository of community-submitted and platform-extracted multi-chain address labels. Each entry includes fields such as name, address, blockchain, source, contributor, etc. We utilize the name field in the table, which contains the naming tags, then apply a set of regular-expression-based rules over the name field to assign coarse-grained functional roles. These rules approximate the aforementioned financial behavioural classes. Table V summarizes the pattern-matching logic used to infer each role: In total, the sampled dataset contains 10,055 transactions between 2023-09-18 and 2025-04-23, covering 815 unique addresses observed as senders or recipients. Using this labeling scheme, we label the address set into four classes: 520 Trader, 33 Bot, 44 Treasury, and 218 Other. These labels serve as training targets for our predictive modeling pipeline, also allowing us to evaluate role inference performance under supervision.
@@
TABLE V: Regular expression used for coarse address role labeling.
@@
V-B Address Representation Learning and Role Inference
U.S. Treasury token transactions exhibit latent financial hierarchy: central actor nodes (e.g., treasuries, issuers, custodians) initiate and coordinate flows, while peripheral nodes (e.g., traders, bots) interact with the core. The negative curvature of hyperbolic space provides a proper inductive bias for embedding such structure (demonstrated in prior work [27]), enabling representations that naturally separate high-degree core nodes near the origin from low-degree boundary nodes along exponentially expanding geodesics. Hence, nodes residing at small hyperbolic radii embody high hierarchy, systemically close to the coordinate origin, possibly multisig treasuries, protocol routers, or liquidity hubs, whereas nodes with larger radii reflect lower-hierarchy participants in the transaction graph. A formal derivation showing that hyperbolic radius encodes hierarchical depth in tree-like graphs is given in Section V-B2, with illustrative support from Figure 3 and an empirical example in Figure 4.
Therefore, we propose a hyperbolic node-level representation learning method combined with a feedforward neural network that integrates transactional features, metadata-driven features including Liquidity-to-Average Ratio (LAR), and hyperbolic (Poincaré) geometry embeddings. The pipeline first encodes latent hierarchy via Poincaré distance-based optimization, then augments each node with hierarchical depth statistics and optional topological features, processed by a neural network classifier to infer address roles.
V-B1 Poincaré Node Representation Learning
Each data point in our framework corresponds to a node $v\in V$ within a token transaction graph $G=(V,E)$, where nodes represent blockchain addresses and directed edges $(u,v)\in E$ denote value transfers. The objective is to learn a high-dimensional representation (embedding vector) $\mathbf{z}_{v}\in\mathbb{B}^{d}$ for each address $v$, where $\mathbb{B}^{d}$ is the $d$-dimensional Poincaré ball: a Riemannian manifold of constant negative curvature:
@@ V-B1 Poincaré Node Representation Learning
where $\|\cdot\|$ is the Euclidean norm. This metric strongly separates nodes at different depths of the hierarchy, with distances growing rapidly near the boundary. Möbius addition is calculated when performing updates while preserving manifold structure, a closed-form generalization of vector translation compatible with hyperbolic geometry:
@@
Each update step is followed by a projection back onto the open ball to ensure validity:
@@
Figure 3: Illustrative example of a blockchain transaction graph and its corresponding Poincaré disk embedding. Nodes emulate three‐tier addresses: the red address anchors global liquidity, the green addresses relay funds and manage liquidity pools, and a ring of peripheral trader blue addresses engage sporadically with the core. Embedding the same adjacency structure in the negatively curved Poincaré model separates tiers by geodesic radius.
@@
where $\gamma>0$ is a margin hyperparameter and $\cdot]_{+}=\max(0,\cdot)$ denotes the hinge operator. To align learned radii with graph-theoretic centrality, we regularize embedding norms against normalized degree:
@@
where $\beta$ (default 0.1) controls the strength of radial regularization. The objective is minimized over epochs using intrinsic gradients computed in the Riemannian manifold, followed by Möbius updates and projection. After optimization, each node $v$ is assigned a $\mathbf{z}_{v}\in\mathbb{B}^{64}$, a 64-dimensional hyperbolic embedding that encodes its topological position and connectivity within the latent hierarchy of the transaction graph.
Refinement via Liquidity-to-Average Ratio (LAR). While Poincaré embeddings encode topological position in the latent transactional hierarchy, other features are related to token flow transferred within transactions, i.e., the temporal liquidity behaviour that distinguishes roles such as passive holders, programmatic bots, or asymmetric treasuries. To incorporate such information, we define the following over nodes computed on transaction flows.
Definition: Liquidity-to-Average Ratio (LAR): Let $(u,v)\in E$ denote a directed transaction edge, and let $\mu_{uv}$ and $\sigma_{uv}$ denote the mean and standard deviation of transfer values from $u$ to $v$ within a time window $t,t+\Delta]$. Let $\texttt{in}(v)$ and $\texttt{out}(v)$ denote the total inflow and outflow of $v$ over the same interval. Then the Liquidity-to-Average Ratio (LAR) for $(u,v)$ is defined as:
@@
Observation. Nodes with low liquidity and irregular behaviour are downweighted during smoothing, preserving geometric sharpness for high-confidence addresses (e.g., well-funded treasuries), while dampening noise from sparsely observed actors.
V-B2 Hyperbolic Radius as a Proxy for Latent Hierarchy
$r_{h}:=\|\phi(u)\|=\tanh\left(\frac{h\cdot\ell}{2}\right),$
where $\ell>0$ is the fixed geodesic step length in the hyperbolic metric. The detailed derivation and proof of the corresponding lemma are provided in Appendix C.
To empirically illustrate that Poincaré embeddings encode latent hierarchy, we visualize address embeddings in the Poincaré disk, with nodes colored by annotated roles and average hyperbolic radius used as a proxy for structural depth. Figure 4 shows three hundred randomly sampled addresses from the RWA Treasury dataset, visualized in two-dimensional hyperbolic space, demonstrating how hyperbolic radius captures latent hierarchical structure.
V-B3 Hierarchical Radius-based Features
To capture the local structure around node $v$, we define its $k$-hop neighbourhood $\mathcal{N}_{k}(v)\subseteq V$ (excluding $v$) and collect neighbour radii $\{r_{u}:u\in\mathcal{N}_{k}(v)\}$. From this, we compute an 11-dimensional hierarchical feature vector detailed in Appendix B.
Figure 4: Illustration of PoincaVec embedding of addresses colored by labels from our U.S. Treasury dataset. Smaller hyperbolic radius indicates higher hierarchy (closer to the origin). Roles like Treasury and Bot appear near the origin, while Trader and Other are peripheral, near the edge.
V-B4 Neural Network Inferencer
After obtaining the Poincaré position embedding and hierarchical radius-based features, we train a supervised neural classifier to infer coarse address roles. Each node $v\in V$ is represented by a concatenated input vector:
  • $\mathbf{z}_{v}\in\mathbb{R}^{d}$ (constrained to $\mathbb{B}^{d}$): Poincaré embedding (cf. Section V-B1);
  • $\mathbf{h}_{v}\in\mathbb{R}^{11}$: hierarchical features computed from local radius-based statistics (cf. Section V-B3);
@@ V-B4 Neural Network Inferencer
V-B5 Experimental Results
We evaluate our role classifier using a stratified train/test split with an 80/20 ratio over the set of labeled addresses with the four coarse-grained role classes (cf. Section V-A). The model is optimized using the AdamW [28] optimizer with a learning rate of $10^{-3}$ and a weight decay of $0.2$. Training is conducted for up to 2k epochs with early stopping based on macro-F1 score on the validation set, using a patience threshold of 10 epochs.
The best-performing model checkpoint (by macro-F1) is selected for final evaluation. We conduct an ablation study using the input features of Poincaré embeddings $\mathbf{z}_{v}\in\mathbb{R}^{64}$, hierarchical descriptors $\mathbf{h}_{v}\in\mathbb{R}^{11}$ ($k=1$), and topology-aware node embeddings $\mathbf{r}_{v}\in\mathbb{R}^{64}$, with feature subsets in combination under various ablation settings.
We compare our model PoincaVec against three established node representation baselines: Node2Vec [21], Role2Vec [1], and FeatherNode [38], included for their demonstrated performance on blockchain transaction graphs in prior evaluations [19, 10]. We adopt the same hyperparameter settings as in prior blockchain address representation learning works [35, 10], setting the context size $10$, embedding dimension $64$, walk length $5$, number of walks per node $10$ for both Node2Vec and Role2Vec, while FeatherNode is configured Singular Value Decomposition (SVD) iterations 20. The baseline Node2Vec [21] captures basic neighbour homophily-based proximity via biased random walks, while Role2Vec [1] models structural equivalence by clustering nodes based on topological statistics into roles (types) and learning role-level embeddings where the nodes share the same role obtain the same embedding vector. FeatherNode [38] incorporates spectral and node-level distributional features from attributed graphs by applying Fourier transforms of empirical distributions to neighbourhood feature aggregations, subsequently compressed using truncated SVD, yielding node embeddings invariant to permutation and sensitive to local graph structure.
Table VI presents the classification performance of PoincaVec and several baseline embedding methods on the RWA U.S. Treasury dataset. PoincaVec (w/ H, w/o T) outperforms all baselines with an F1 of 0.693. Integrating hierarchical radius features with topology-aware node vectors derived from truncated random walks (w/ H, w/ T) yields an F1 score of 0.726. These results demonstrate that the PoincaVec embeddings, produced by the proposed method enriched with structural features, provide competitive representational capacity for role inference in blockchain transaction graphs.
TABLE VI: Role classification performance on the RWA U.S. Treasury dataset. H: hierarchical radius-based features; T: topology-aware node features.
@@ V-B5 Experimental Results
V-B6 Extended Evaluation on Public Blockchain Transaction Datasets
To further evaluate the generalization capacity of our PoincaVec architecture beyond the RWA role classification task, we test it on multiple publicly available blockchain transaction graph datasets with labels. Detailed statistics and descriptions of the datasets are in Appendix A. These datasets include addresses labeled as fraudulent or scam-related, whose topological irregularities render them particularly amenable to detection using curvature-aware embeddings and hierarchical structural descriptors. Given that fraud or anomalous addresses often lie structurally at the periphery or occupy non-homophilic positions in the graph, the curvature-aware inductive bias of PoincaVec may help reveal subtle hierarchical distinctions absent in flat Euclidean spaces.
TABLE VII: Role classification performance across four blockchain transaction datasets.